OpenStudy (anonymous):

Solve 5x^2 = -30x - 65.

OpenStudy (anonymous):

I think it's x= -3 + 4i but i'm not sure.

OpenStudy (campbell_st):

well rewrite it and divide every term by 5 \[x^2 + 6x = 13\] complete the square \[x^2 + 6x + 9 = 13 + 9\] which becomes \[(x + 3)^2 = 22\] now you can solve for x

OpenStudy (anonymous):

I'm not sure what to substitute in for x... @campbell_st

OpenStudy (anonymous):

Wait, is it x = -6 + 2i

OpenStudy (anonymous):

@campbell_st

OpenStudy (phi):

How were you taught to solve this type problem?

OpenStudy (anonymous):

I wasn't that's the thing.

OpenStudy (phi):

There are 3 ways (1) factoring (won't work for this problem) (2) quadratic formula (3) complete the square any sound familiar?

OpenStudy (anonymous):

I know factoring..that's it

OpenStudy (phi):

I would start with 5x^2 = -30x - 65 and put it into standard form. Can you do that ?

OpenStudy (campbell_st):

ok so if \[(x + 3)^2 = 22\] taking the square root of both sides gives \[x + 3 = \pm \sqrt{22}\] perhaps this helps... to solve for x

OpenStudy (phi):

standard form means ax^2 + bx + c = 0